A block of mass 1kg is suspended freely by a thread. A bullet of mass 3kg is fired horizontally at the block and it becomes embedded in it. The block swings to one side using a vertical distance of 50cm. With what speed did the bullet hit the block?


1
Expert's answer
2021-08-04T16:24:22-0400

Given:

"m=1\:\rm kg"

"M=3\:\rm kg"

"h=0.5\:\rm m"

Let the initial speed of a bullet is denoted as "u". The law of conservation of momentum gives

"Mu=(m+M)v"

The law of conservation of energy gives

"\frac{(M+m)v^2}{2}=(M+m)gh"

Hence, we get

"u=\frac{M+m}{M}v=\frac{M+m}{M}\sqrt{2gh}"

"u=\frac{3+1}{3}\sqrt{2*9.8*0.5}=4.17\:\rm m/s"